OpenStudy (anonymous):

Write the expression 2i/3-4i as a complex number in standard form.

jhonyy9 (jhonyy9):

do you know what is the conjugate of denominator ?

OpenStudy (anonymous):

2\[\frac{ 2i }{ 3-4i }*\frac{ 3+4i }{ 3+4i }\]

OpenStudy (anonymous):

\[\frac{ 6i-8 }{25 }\]